# LLM-as-a-qualitative-judge: automating error analysis in natural language generation

### v2

@@ 4.2 Per-instance analysis
[…]
@@
Table 3: Performance of various approaches and LLMs in issue clustering. Results averaged over 3 runs from different random seeds. Agreement in cluster assignment measured using Adjusted Rank Index (ARI) and in cluster descriptions using LLM-judged Semantic Label Consistency (SLC, evaluated by Claude-3.7-Sonnet). Subscripts ${}_{\text{real}}$ and ${}_{\text{syn}}$ indicate tests on the real and synthetic data respectively. “N/a” indicates the metric is not applicable since classic approaches do not generate cluster names.
@@ 4.3 Issue clustering
Figure 4 demonstrates examples of confusion matrices for several datasets. Pronounced diagonals and matching cluster names illustrate the strong capabilities of LLM-as-a-qualitative-judge to output comprehensive issue types reports. Due to the inherent subjectivity of clustering task, we observe occasional merging or splitting of annotator’s clusters, e.g. clusters “Wrong topping” and “Wrong variables” were merged by LLM-as-a-qualitative-judge into one cluster “Entity Mislabeling” for the Pizza ordering dataset.
Table 4 demonstrates the non-uniformity of LLM-as-a-qualitative-judge’s performance across datasets, i.e. issue type reports generated by LLM-as-a-qualitative-judge reach high agreement with annotator’s reports for certain datasets and rather low agreement for other datasets. We attribute this effect to the imperfections of per-instance analysis discussed above (which we believe could be further alleviated in future work) and the inherent subjectivity of the issue explanation and clustering tasks. Appendix E present the examples of generated error type reports, per-instance issue explanations, and confusion matrices for all considered datasets. We also provide an archive with experimental results in the project repository [^4] A zip archive with experimental results is available at https://github.com/tunde-ajayi/llm-as-a-qualitative-judge/tree/main/data.
@@
5 Related work
Quantitative LLM-based evaluation.
While using commercial LLMs for evaluation remains common practice, one line of work (Kim et al., 2024b, c; Pombal et al., 2025) focuses on tuning open-source LLMs on the synthetically generated evaluation data, to ensure reproducibility of evaluation. Other works improve quantitative LLM-as-a-judge by conducting more fine-grained evaluation, e.g. using multiple evaluation criteria (Liang et al., 2023; Fu et al., 2024) or selecting evaluation criteria individually per instance (Ye et al., 2024; Cook et al., 2024; Kim et al., 2025). Composite evaluation approaches such as FactScore (Min et al., 2023) or RAGChecker (Ru et al., 2024) use LLMs in the intermediate evaluation steps.
Qualitative LLM-based evaluation.
LLM-generated qualitative error explanations are often used to improve the precision of quantitative evaluation (Zeng et al., 2024; Ye et al., 2024) or to explain the assigned quantitative scores to a developer (Xu et al., 2023; Jiang et al., 2024). Such approaches only output per-instance explanations, and a substantial human effort is still needed to read all of them. Certain works (Jiang et al., 2024; Kasner et al., 2024; Perrella et al., 2022) focus on outputting aggregated reports of frequent errors, but with a (limited) predefined error set, i.e. they solve the task of error classification. In contrast to these efforts, LLM-as-a-qualitative-judge outputs an aggregated report of issue types discovered in an open-ended manner, i.e. without any predefined issue set.
Meta-evaluation.
A line of community efforts (Zeng et al., 2024; Lambert et al., 2025; Hada et al., 2024; Bavaresco et al., 2024) is devoted to an important task of meta-evaluating LLM-as-a-judge, i.e. collecting human annotations for various tasks, domains, or languages, and evaluating how closely LLMs mirror human judgments. Certain task-specific datasets (Freitag et al., 2021) can be used to meta-evaluate fine-grained issue detection. Our work further contributes to this direction by the release of a meta-evaluation dataset, containing qualitative issue explanations for 12 datasets from 7 tasks and their per-dataset clustering.
Clustering with LLMs.
Earlier works (Petukhova et al., 2024; Miller and Alexander, 2024) demonstrate advantages of leveraging LLM-derived embeddings in place of traditional TF-IDF or BERT vectors in standard clustering algorithms. More recent works employ LLMs directly to cluster textual data. Viswanathan et al. (2024) instruct a GPT-3.5 model to cluster the provided data given few-shot demonstrations. Huang and He (2024) transform clustering into a two-stage classification task: first prompting an LLM to infer a set of candidate clusters for the dataset, then prompting it to assign the best cluster to each instance. ClusterLLM Zhang et al. (2023) uses an instruction-tuned LLM to guide clustering, i.e., to decide which clusters to merge. In our work, we propose an alternative intuitive approach for LLM-based clustering. Our approach can also be extended in the future with the listed strategies.
6 Conclusion
In this work, we present LLM-as-a-qualitative-judge, a novel approach for generating structured reports summarizing key types of issues in a given NLG system. We hope that this approach will help developers to spot more easily issues and artifacts in their NLG systems. For example, from reports shown in Figure 3, a developer may decide to revisit the training data mixture (left) or a task metric (right). The LLM-as-a-qualitative-judge approach can also be easily extended to compare two NLG systems, by using two separate issue counters (one per dataset) in the cumulative clustering algorithm. Our code and data are available at https://github.com/tunde-ajayi/llm-as-a-qualitative-judge.
Acknowledgments
Appendix A Prompts
@@
Appendix B Clustering experiment setup
Appendix C Models
Model BibTeX License Model Repository
GPT-4o OpenAI et al. (2024) Proprietary https://platform.openai.com/docs/models/gpt-4o
Gemini-2.0-Flash Anil et al. (2023) Proprietary https://deepmind.google/technologies/gemini/flash/
Qwen-2.5 Qwen et al. (2025) Apache 2.0 https://huggingface.co/collections/Qwen/qwen25-66e81a666513e518adb90d9e
DeepSeek-R1-Distill-Llama-8B DeepSeek-AI (2025) Llama 3.1 Community License https://huggingface.co/deepseek-ai/DeepSeek-R1-Distill-Llama-8B
Aya-Expanse-8B Dang et al. (2024) Creative Commons Attribution Non Commercial 4.0 https://huggingface.co/CohereLabs/aya-expanse-8b
Llama-3.1-8B-Instruct Grattafiori et al. (2024) Llama 3.1 Community License https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct
Llama-3.2-1B-Instruct Grattafiori et al. (2024) Llama 3.2 Community License https://huggingface.co/meta-llama/Llama-3.2-1B-Instruct
Ministral-8B-Instruct Mistral AI (2024) Mistral AI Research License https://huggingface.co/mistralai/Ministral-8B-Instruct-2410
Solar-10.7B Kim et al. (2024a) Creative Commons Attribution Non Commercial 4.0 https://huggingface.co/upstage/SOLAR-10.7B-Instruct-v1.0
Vicuna-1.5-13B Chiang et al. (2023) Llama 2 Community License Agreement https://huggingface.co/lmsys/vicuna-13b-v1.5
Command-R-35B Gomez and for AI (2024) Creative Commons Attribution Non Commercial 4.0 https://huggingface.co/CohereLabs/c4ai-command-r-v01
Table 5: References to the used LLMs; all LLMs allow use for research.
Appendix D Datasets
Dataset name Dataset reference License
Natural Language Generation
FLASK data mix: Self-Instruct, WizardLM, Koala, CommonSense QA Wang et al. (2023); Xu et al. (2024); Geng et al. (2023); Talmor et al. (2021) Apache 2.0, MIT, Apache 2.0, Creative Commons Attribution 4.0
WMT’22 Kocmi et al. (2022) Apache 2.0
Elitr-Bench Thonet et al. (2025) Attribution 4.0 International
PIZZA Arkoudas et al. (2022) Attribution-NonCommercial 4.0 International
GSM8K Cobbe et al. (2021) MIT License
ParaDetox Dementieva et al. (2024) OpenRAIL++
Retrieval-augmented QA
MKQA (ru) Longpre et al. (2021) Creative Commons Attribution-ShareAlike 3.0 Unported License
BioASQ Krithara et al. (2023) Attribution 2.5 Generic
RobustQA Santhanam et al. (2022); Han et al. (2024, 2023) Apache-2.0
SearchQA Dunn et al. (2017) BSD 3-Clause
SyllabusQA Fernandez et al. (2024) Attribution-NonCommercialShareAlike
Appendix E Examples

### v4

@@ 5.3 Examples of per-issue analysis
$_{\text{real}} ARI$$_{\text{syn}}$ $_{\text{real}} SLC$$_{\text{syn}}$
@@
Table 3: Performance of various approaches and LLMs in issue clustering. Results averaged over 3 runs from different random seeds. Agreement in cluster assignment measured using Adjusted Rank Index (ARI) and in cluster descriptions using LLM-judged Semantic Label Consistency (SLC). Subscripts $_{\text{real}}$ and $_{\text{syn}}$ indicate tests on the real and synthetic data respectively. “N/a” indicates the metric is not applicable since classic approaches do not generate cluster names.
The second occasional problem in per-instance analysis is the oversimplification of an issue, especially for more unexpected issues, such as an ambiguous task input or an error in evaluation. In the example in row 5, the issue is an ambiguous user question, i.e., both the ground truth and the generated response are correct and provide two different interpretations of the user question. However, GPT-4o and Qwen-2.5-7B report the over-simplified issue of the generated response not providing the same answer as the ground truth response.
Finally, the third occasional reason for a per-instance issue explanation not being accepted by the evaluator LLM, is the subjectivity of some issues in the dataset. For example, a human-annotated issue in row 6, “The generation was stopped too early because of the reached maximum new tokens limit”, is evaluated to be not equivalent to the LLM-as-a-qualitative-judge-generated issue “The generated response provides an incomplete overview <…>”. While these two issue explanations are indeed different, they are both correct, and the LLM-generated explanation follows from the annotator’s explanation.
To alleviate potential negative effects from erroneous issue explanations, we recommend developers to check a couple of examples of each issue, which are output by LLM-as-a-qualitative-judge in addition to the issue names and descriptions.
@@ 5.4 Issue clustering
Figure 5 demonstrates examples of confusion matrices for several datasets. Pronounced diagonals and matching cluster names illustrate the strong capabilities of LLM-as-a-qualitative-judge to output issue types reports that resemble the issue reports produced by humans. Due to the inherent subjectivity of clustering task, we observe occasional merging or splitting of annotator’s clusters, e.g. clusters “Wrong topping” and “Wrong variables” were merged by LLM-as-a-qualitative-judge into one cluster “Entity Mislabeling” for the Pizza ordering dataset.
To sum up, our results demonstrate the effectiveness of the proposed cumulative clustering approach to produce issue reports that resemble the ones produced by humans. For practical applications, we recommend using recent models such as GPT-4o or Gemini-2.5-Flash.
6 Discussion
In this section, we discuss potential extensions of the proposed approach.
Issues prefiltering.
As discussed in Section 2, LLM-as-a-qualitative-judge operates only over instances which received low scores from a quantitative task metric. Such prefiltering could in principle be removed and incorporated directly into per-instance analysis by modifying its prompt, e.g. “Output what is an issue with this instance. If there is no issue, output ’No issue’”. Instances with predictions “No issue” then would be discarded from issue clustering. However, in preliminary experiments we found that LLM are prone to making up issues for fully correct instances. Hence, we do not recommend removing the prefiltering step (at least with the current state of LLMs), which is a reasonable design since LLM-as-a-qualitative-judge is an error analysis method.
Multiple issues per instance.
LLM-as-a-qualitative-judge could be easily extended to detect multiple issues per instance, by modifying the prompt used for per-instance analysis and going through the generated issues one-by-one in the issue clustering step. However, same as with a previous discussion point, in our preliminary experiments we found that LLMs are prone to generating non-existing issues in such a scenario. For example, GPT-4o tends to generate a constant number of issues for any instance (in particular, 3). In practice, we believe that our design with one issue per instance is reasonable, since most of the erroneous instances have only one issue. Furthermore, even if some instances have repeating issues, our algorithm would still capture most of the issues in the dataset due to issue repetition.
Pairwise comparison of models.
LLM-as-a-qualitative-judge can be straightforwardly extended to perform pairwise comparison of models, by running per-instance analysis for outputs of both models and using two counters (one per model) in the issue clustering step.
Use without ground truth labels.
LLM-as-a-qualitative-judge can be straightforwardly used without ground truth labels, i.e. as an unsupervised evaluation metric, if the LLM internal knowledge is sufficient to understand errors in a given task. LLM-as-a-qualitative-judge can also be provided with additional evaluation metadata, e.g. score rubrics used in quantitative evaluation.
7 Related work
Quantitative LLM-based evaluation.
While using commercial LLMs for evaluation remains common practice, one line of work (prometheus; prometheus2; multiprometheus) focuses on tuning open-source LLMs on the synthetically generated evaluation data, to ensure reproducibility of evaluation. Other works improve quantitative LLM-as-a-judge by conducting more fine-grained evaluation, e.g. using multiple evaluation criteria (liang2023holistic; fu-etal-2024-gptscore) or selecting evaluation criteria individually per instance (flask; checklists; biggen). Composite evaluation approaches such as FactScore (factscore) or RAGChecker (ragchecker) use LLMs in the intermediate evaluation steps.
Qualitative LLM-based evaluation.
LLM-generated qualitative error explanations are often used to improve the precision of quantitative evaluation (zeng2024evaluating; flask) or to explain the assigned quantitative scores to a developer (instructscore; tigerscore). Such approaches only output per-instance explanations, and a substantial human effort is still needed to read all of them. Certain works (tigerscore; factgenie; matese) focus on outputting aggregated reports of frequent errors, but with a (limited) predefined error set, i.e. they solve the task of error classification. In contrast to these efforts, LLM-as-a-qualitative-judge outputs an aggregated report of issue types discovered in an open-ended manner, i.e. without any predefined issue set.
Meta-evaluation.
A line of community efforts (zeng2024evaluating; lambert-etal-2025-rewardbench; metal; bavaresco2024llmsinsteadhumanjudges) is devoted to an important task of meta-evaluating LLM-as-a-judge, i.e. collecting human annotations for various tasks, domains, or languages, and evaluating how closely LLMs mirror human judgments. Certain task-specific datasets (freitag-etal-2021-experts) can be used to meta-evaluate fine-grained issue detection. Our work further contributes to this direction by the release of a meta-evaluation dataset, containing qualitative issue explanations for 12 datasets from 7 tasks and their per-dataset clustering.
Clustering with LLMs.
Earlier works (DBLP:journals/corr/abs-2403-15112; DBLP:journals/corr/abs-2405-07278) demonstrate advantages of leveraging LLM-derived embeddings in place of traditional TF-IDF or BERT vectors in standard clustering algorithms. More recent works employ LLMs directly to cluster textual data. DBLP:journals/tacl/0002GGLWN24 instruct a GPT-3.5 model to cluster the provided data given few-shot demonstrations. DBLP:journals/corr/abs-2410-00927 transform clustering into a two-stage classification task: first prompting an LLM to infer a set of candidate clusters for the dataset, then prompting it to assign the best cluster to each instance. ClusterLLM DBLP:conf/emnlp/0001WS23 uses an instruction-tuned LLM to guide clustering, i.e., to decide which clusters to merge. In our work, we propose an alternative intuitive approach for LLM-based clustering. Our approach can also be extended in the future with the listed strategies.
8 Conclusion
In this work, we present LLM-as-a-qualitative-judge, a novel approach for generating structured reports summarizing key types of issues in a given NLG system. We hope that this approach will help developers to spot more easily issues and artifacts in their NLG systems.
Limitations
As any LLM-based system, LLM-as-a-qualitative-judge can make occasional mistakes in analysis or clustering. In Section 5, we discuss types of such mistakes and recommend checking a couple of examples of each issue, which are also output by LLM-as-a-qualitative-judge.
Regarding limitations of the evaluation methodology, despite our efforts in considering a diverse set of tasks, domains, and LLMs, we acknowledge the infeasibility of covering the entire breadth of NLG applications and models in our study. Another limitation is that we mainly focus on English. We believe our findings will transfer to other languages, with the use of strong recent multilingual LLMs, but acknowledge that the reliability of LLM-as-a-qualitative-judge in multilingual studies requires a separate study.
Broader impact
We acknowledge that as any LLM-based system, LLM-as-a-qualitative-judge can make errors which could propagate to the downstream systems and decrease their performance. For example, if developers rely solely on the issue names formulated by Judge, this could occasionally lead to unnecessary or even harmful modifications of their NLG systems. This could also happen in case of misinterpetation of an issue by a developer due to issue subjectivity. To reduce such risks, we recommend developers to check examples of issue types, which are also output by LLM-as-a-qualitative-judge, in addition to the issue names and description.
Acknowledgments
A part of this work was carried out within the framework of the AugmentIA Chair, supported by the Fondation Grenoble INP thanks to the patronage of Artelia Group, and is affiliated with Laboratory of informatics in Grenoble (LIG). A part of this work received government funding managed by the French National Research Agency under France 2030, reference ANR-23-IACL-0006.
Appendix A Details on data annotation
Per-instance analysis.
The core of our meta-evaluation strategy is to collect manual annotations of failure cases for a set of instances from various tasks and domains. For each instance, consisting of a task input, a ground truth response, a generated response, a description of a task metric, and optionally retrieved documents, an annotator’s task is to formulate what the particular issue is in this instance. We then prompt an evaluator LLM to judge whether the issue explanation determined by the LLM-as-a-qualitative-judge for a particular instance matches the issue determined by the human annotator. The outputs from the evaluator LLM are binary and can be accumulated into a per-instance analysis accuracy score.
Issue clustering.
Dataset composition.
For each considered dataset, we manually label failures in up to 40 generations from one of the open-source LLMs (Qwen-2.5-1b, Llama-3.2-1b, Command-R-35b, Vicuna-1.5-13B).
Annotation details.
All the labeling was performed by the authors of the paper in Google Spreadsheets [^4] https://docs.google.com/spreadsheets. Each instance was annotated by one author. Authors of the paper are PhD students in the NLP field or have already completed their PhD in NLP and are employed as NLP researchers.
Time needed for data annotation varies between tasks: it took us 1—6 hours per task. Tasks with longer inputs, e.g. RA-QA, and from more complex domains, e.g. biomedical, take more time to annotate, e.g. they require reading the retrieved documents carefully.
Below is an annotation instruction:
For each example, consisting of a user prompt, a ground truth label, an LLM generation, and optionally retrieved documents, an annotator’s task is to formulate what is a particular failure case in this example. Identify only one, most important specific, clearly visible issue in each test case. Please formulate the detected issue as a clear, full sentence, e.g. "The generated response is in German instead of French which is the language of the user input" or "The retrieved documents are from a datastore which is irrelevant to the given user question". Please add your annotations in the following Google Spreadsheet: [link], column “Per-instance analysis”. You can skip instances (rows) for which you feel that you do not have enough expertise to detect an issue, which have multiple issues, or for which ground truth labels appear to be wrong. After annotating per-instance analysis, suggest a clustering of the detected issues, i.e. how would you group them, and add the corresponding cluster indexes and names in columns “Cluster index” and ‘‘Cluster name’’.
Inter-annotator agreement.
We measure the inter-annotator agreement on a subset of 100 instances, i.e. each of these instances was labeled by two annotators and then we computed their agreement using the same evaluator LLM as in other experiments, i.e. Claude-3.7-sonnet. The resulting inter-annotator agreement was 57% (percentage of cases when two annotators suggested the same issue, as judged by Claude-3.7-sonnet), i.e. the similar range as the scores we obtain in Table 2.
The main factor contributing to the moderate agreement is the subjectivity of issue analysis. For example, in a situation when generation was stopped due to reaching the maximum new tokens limit, one annotator said “The response is incomplete” and another annotator said “Generation was stopped too early”. Both denote the same root issue, but are formulated differently and Claude judges these comments as different.
Appendix B Further details on the experimental setup
Case study.
For each of the three considered BigBenchHars tasks, we build a simple initial generative pipeline. This pipeline is then improved in two rounds by generating issue reports with LLM-as-a-qualitative-judge. Configurations of the initial pipeline are as follows. System prompt: “You are a helpful assistant. Output your answer after a final separator ‘Answer:’”. Generation hyperparameters: all hyperparameters set to default values from the HuggingFace or OpenAI API, plus setting maximum new tokens or 500 for HuggingFace models. The final answers are obtained by cropping the content after a final separator “Answer:” and applying a .strip() python function. Evaluation function: exact match with ground truth. LLM-as-a-qualitative-judge is run with GPT-4.1 and providing a one-sentence description of a task metric, i.e. “Evaluation is conducted using exact matching between the ground-truth label and the content of the generated response after the final separator ‘Answer:”’.
Meta-evaluation experiments.
For each instance, LLM-as-a-qualitative-judge is provided with a task input, a ground truth response, a generated response, 5 retrieved documents (only for RA-QA), and a short task comment. The task comment describes the task metric (in one sentence), provides a comment on the nature of ground truth responses (either that it is the expected answer or that it is only one of the possible correct answers), and also contains a comment that retrieval-augmented generation (RAG) or Chain-of-Thought (COT) prompting was used, when applicable (6 datasets with RAG and 2 datasets with COT). The used task metrics are binary LLM-as-a-judge (the generated response is accepted or not) or binary Match (outputs True is one of the ground truth answers is contained a substring in the generated response, and False otherwise). Task comments for all datasets are also presented in Appendix H.
For LLM-as-a-qualitative-judge, open-source LLMs are run on a single V100 GPU with greedy decoding ($\sim$20 GPU-hours in total). Commercial LLMs are run via API with requesting json output format.
The time of running the LLM-as-a-qualitative-judge algorithm depends on the setting (commercial vs open-source LLMs, type of GPU etc) and in our experiments was taking 2—30 min, i.e. reasonably short on the scale of the time needed to develop an NLG system.
Meta-evaluation of an evaluator LLM.
To ensure the reliability of the evaluator LLM, we collected a small meta-evaluation dataset of 50 instances from 4 datasets (MKQA (ru), RobustQA Writing, FLASK, MultiDetox), where the equivalence of the LLM-as-a-qualitative-judge’s and human annotator’s per-instance analysis was judged by a human annotator and can be compared to the evaluator LLM’s verdict. Strong commercial LLMs, such as GPT-4o, Gemini-2.0-Flash, and Claude-3.7-Sonnet, achieved a meta-evaluation accuracy of 85-90% on this dataset, and an open-source Solar-10.7B (kim2023solar) achieved a meta-accuracy of 60%. In all the experiments, we use claude-3-7-sonnet-20250219 as the evaluator LLM, to avoid using the same LLM for analysis and for evaluation.
Appendix C Clustering experiment setup
Appendix D Models
Model BibTeX License Model Repository
GPT-4o openai2024gpt4technicalreport Proprietary https://platform.openai.com/docs/models/gpt-4o
Gemini-2.0-Flash geminiteam2025geminifamilyhighlycapable Proprietary https://deepmind.google/technologies/gemini/flash/
Qwen-2.5 qwen2025qwen25technicalreport Apache 2.0 https://huggingface.co/collections/Qwen/qwen25-66e81a666513e518adb90d9e
DeepSeek-R1-Distill-Llama-8B deepseekai2025deepseekr1incentivizingreasoningcapability Llama 3.1 Community License https://huggingface.co/deepseek-ai/DeepSeek-R1-Distill-Llama-8B
Aya-Expanse-8B dang2024ayaexpansecombiningresearch Creative Commons Attribution Non Commercial 4.0 https://huggingface.co/CohereLabs/aya-expanse-8b
Llama-3.1-8B-Instruct grattafiori2024llama3herdmodels Llama 3.1 Community License https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct
Llama-3.2-1B-Instruct grattafiori2024llama3herdmodels Llama 3.2 Community License https://huggingface.co/meta-llama/Llama-3.2-1B-Instruct
Ministral-8B-Instruct MistralAI2024Ministraux Mistral AI Research License https://huggingface.co/mistralai/Ministral-8B-Instruct-2410
Solar-10.7B kim2023solar Creative Commons Attribution Non Commercial 4.0 https://huggingface.co/upstage/SOLAR-10.7B-Instruct-v1.0
Vicuna-1.5-13B vicuna2023 Llama 2 Community License Agreement https://huggingface.co/lmsys/vicuna-13b-v1.5
Command-R-35B commandr Creative Commons Attribution Non Commercial 4.0 https://huggingface.co/CohereLabs/c4ai-command-r-v01
Table 4: References to the used LLMs; all LLMs allow use for research.
Appendix E Datasets
Dataset name Dataset reference License
Natural Language Generation
FLASK data mix: Self-Instruct, WizardLM, Koala, CommonSense QA wang-etal-2023-self-instruct; xu2024wizardlm; koala_blogpost_2023; DBLP:conf/nips/TalmorYBBGCB21 Apache 2.0, MIT, Apache 2.0, Creative Commons Attribution 4.0
WMT’22 kocmi-etal-2022-findings Apache 2.0
Elitr-Bench DBLP:conf/coling/ThonetBR25 Attribution 4.0 International
PIZZA DBLP:journals/corr/abs-2212-00265 Attribution-NonCommercial 4.0 International
GSM8K DBLP:journals/corr/abs-2110-14168 MIT License
ParaDetox dementieva2024overview OpenRAIL++
Retrieval-augmented QA
MKQA (ru) longpre-etal-2021-mkqa Creative Commons Attribution-ShareAlike 3.0 Unported License
BioASQ krithara2023bioasq Attribution 2.5 Generic
RobustQA lotte; DBLP:conf/emnlp/Han00XWLWMC24; Han2023 Apache-2.0
SearchQA DBLP:journals/corr/DunnSHGCC17 BSD 3-Clause
SyllabusQA DBLP:conf/acl/FernandezSL24 Attribution-NonCommercialShareAlike
BigBenchHard
Date understanding; Word sorting; Movie recommendation bbh MIT
Appendix F Per-dataset results
Table 6 presents per-dataset results for for GPT-4o as LLM-as-a-qualitative-judge.
@@ Appendix F Per-dataset results
Appendix G Prompts
@@
Appendix H Examples